# On the Real Zeroes of Half-integral Weight Hecke Cusp Forms, II

### v1

@@ 9. Relating the Fourth moment to a random model
Applying Poisson summation to the $m_{1}$-sum gives
$\frac{1}{[c,d]}\sum_{v}e\left(\frac{xv}{[c,d]}\right)\int\limits_{\mathbb{R}}y^{-3/4}e\left(\frac{2\sqrt{ym_{2}u}}{c}-\frac{yv}{[c,d]}\right)W_{K}^{(2)}\left(\frac{yj}{|d|},\frac{m_{2}j}{|d|},\frac{K^{2}c}{8\pi\sqrt{ym_{2}u}}\right)\,\mathrm{d}y.$
Similarly applying Poisson summation to the $m_{2}$-sum gives that after some computations our double sum equals
$\sum_{x\,([c,d])}\sum_{w\,([c,d])}S\left(xw,u;c\right)\chi_{d}(x)\chi_{d}(w)\frac{1}{[c,d]^{2}}\sum_{v}\sum_{\eta}e\left(\frac{xv+w\eta}{[c,d]}\right)\cdot\mathcal{I},$
@@
We truncate the $y$-integral smoothly to $|y|\leq K^{1+\varepsilon}|d|$ with a negligible error and then apply Lemma 5.9 with the choices $Q=V=V_{1}=K|d|/j$, $X=\sqrt{z}d/Kcj$, and $Y=\sqrt{zuK|d|/j}c^{-1}$. It is straightforward to check that the conditions of the lemma are met under (9.3). We conclude that
@@
$\frac{-2\sqrt{2\pi}}{\sqrt{u}}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{(j,d)=1}\frac{1}{j}\textnormal{Im}\bigg(e^{-\pi i/2}\sum_{c=1}^{\infty}\frac{1}{[c,d]^{2}}\sum_{v}\sum_{\eta}\sum_{x\,([c,d])}\sum_{w\,([c,d])}\chi_{d}(x)\chi_{d}(w)$
(9.4) $\times S(xw,u;c)e\left(\frac{xv+w\eta}{[c,d]}\right)\int\limits_{\mathbb{R}}z^{-1}e\left(z\left(\frac{u[c,d]}{c^{2}v}-\frac{\eta}{[c,d]}\right)\right)W_{K}^{(2)}\left(\frac{zu[c,d]^{2}j}{v^{2}c^{2}|d|},\frac{zj}{|d|},\frac{K^{2}c^{2}v}{8\pi zu[c,d]}\right)\,\mathrm{d}z\bigg)+O\left(X^{3/4}K^{1+\varepsilon}\right).$
@@
$v\ll\frac{[c,d]|d|uK^{\varepsilon}}{c^{2}K}$
and the same holds for $\eta$. Note that as $|d|\ll X\ll\sqrt{K}$ this shows that the integral is negligible unless $v,\eta\leq K^{\varepsilon}/c$, which effectively truncates also the $c$-sum at $K^{\varepsilon}$. Note also that we may truncate the $\eta$-sum at $X^{2}$ as otherwise the derivative of the phase function is $\gg[c,d]^{2}/c^{2}v$. Now using the first derivative test and estimating everything else trivially using the Weil bound for Kloosterman sums shows that (9) is $\ll X^{3/4}K^{1+\varepsilon}$.
@@
$-2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}c_{n}(u)\frac{\chi_{d}(u)}{\sqrt{u}}$
@@
where the summand on the right-hand side is a multiplicative function of $c$. We also have by a simple calculation that, for $\textnormal{Re}(x+y)>0$,
$\sum_{c=1}^{\infty}\frac{\varphi(c)}{c^{1+x+y}}\cdot\frac{1}{d}\prod_{\begin{subarray}{c}p|d\\ p|c\end{subarray}}(-p)$ $=\frac{1}{d}\prod_{p}\left(1+\sum_{j=1}^{\infty}\frac{\varphi(p^{j})}{p^{j(1+x+y)}}\prod_{\begin{subarray}{c}q|d\\ q|c\end{subarray}}(-q)\right)$
@@
Making the change of variables $z\mapsto cz$, using the approximation (5.11) and Poisson summation we see that the integral above equals
@@
$=\frac{1}{(2\pi i)^{2}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}\left(\frac{hcuj}{K|d|v^{2}}\right)^{-x}\left(\frac{hcj}{K|d|}\right)^{-y}\hbar_{x+y}\left(\frac{K^{2}|d|v}{8\pi hu}\right)\,\frac{\mathrm{d}x\,\mathrm{d}y}{xy}$
$=\frac{1}{(2\pi i)^{3}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}\left(\frac{hcuj}{K|d|v^{2}}\right)^{-x}\left(\frac{hcj}{K|d|}\right)^{-y}\left(\frac{8\pi hu}{K^{2}|d|v}\right)^{z}\widetilde{\hbar}_{x+y}(z)\,\mathrm{d}z\,\frac{\mathrm{d}x\,\mathrm{d}y}{xy},$
@@
$2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)$
$\qquad\times\sum_{v|u}\frac{1}{(2\pi i)^{3}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\widetilde{\hbar}_{x+y}(z)$
@@
$=2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)\sum_{v|u}$
$\qquad\times\frac{1}{(2\pi i)^{3}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\widetilde{\hbar}_{x+y}(z)$
@@
$2K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)\sum_{v|u}$
$\qquad\times\frac{1}{(2\pi i)^{3}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{0}^{\infty}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\xi^{(x+y)/2-z}$
@@
$2K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)\sum_{v|u}\frac{1}{(2\pi i)^{2}}\int\limits_{(\varepsilon)}\int\limits_{(\varepsilon)}\int\limits_{0}^{\infty}(2\pi K)^{-z}u^{z-x}(8\pi)^{z}e^{x^{2}+(z-x)^{2}}v^{2x-z}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\xi^{-z/2}$
@@
(9.7) $-2\zeta(0)K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\varepsilon)}\int\limits_{0}^{\infty}e^{2x^{2}}u^{-x}v^{2x}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\prod_{p|d}\left(1-p^{-1}\right)\left(\frac{1-p}{p-1}\right)\,\mathrm{d}\xi\frac{\mathrm{d}x}{x^{2}}.$
Simplifying this by recalling that $\zeta(0)=-1/2$ and making a change of variables in the $\xi$-integral we see that our integral in (9) equals, up to a negligible error,
$2\widehat{h}(0)K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(u)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\varepsilon)}e^{2x^{2}}u^{-x}v^{2x}\,\frac{\mathrm{d}x}{x^{2}}.$
@@
$M_{2}(X;d)$ $:=(\log K)\sum_{n\leq K^{4\theta_{J}}}\frac{h_{2}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u)$
@@
$\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{f\in B_{k}}\omega_{f}L\left(\frac{1}{2},f\otimes\chi_{d}\right)^{2}M_{f}(d)^{4}=\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathbb{E}\left(L(X;d,k)M_{2}(X;d)\right)$
(9.8) $+2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\theta_{J}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\frac{\varphi(d)}{d}\phi\left(\frac{|d|}{X}\right)\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\frac{1}{2\pi i}\int\limits_{(\varepsilon)}\frac{e^{2x^{2}}}{x^{2}}\cdot u^{-x}\sum_{v|u}v^{2x}\,\mathrm{d}x+O\left(X^{3/4}K^{1+\varepsilon}\right)$

### v2

@@ 9. Relating the Fourth moment to a random model
Applying Poisson summation to the $m_{1}$-sum gives that it equals
$\frac{1}{[c,d]}\sum_{v\in\mathbb{Z}}e\left(\frac{xv}{[c,d]}\right)\int\limits_{\mathbb{R}}y^{-3/4}e\left(\frac{2\sqrt{ym_{2}u}}{c}-\frac{yv}{[c,d]}\right)W_{K}^{(2)}\left(\frac{yj}{|d|},\frac{m_{2}j}{|d|},\frac{K^{2}c}{8\pi\sqrt{ym_{2}u}}\right)\,\mathrm{d}y.$
Similarly applying Poisson summation to the $m_{2}$-sum gives, after some computations, that our double sum equals
$\sum_{x\,([c,d])}\sum_{w\,([c,d])}S\left(xw,u;c\right)\chi_{d}(x)\chi_{d}(w)\frac{1}{[c,d]^{2}}\sum_{v\in\mathbb{Z}}\sum_{\eta\in\mathbb{Z}}e\left(\frac{xv+w\eta}{[c,d]}\right)\cdot\mathcal{I},$
@@
We truncate the $y$-integral smoothly to $|y|\leq K^{1+\varepsilon}|d|/j$ with a negligible error and then apply Lemma 5.9 with the choices $Q=V=V_{1}=K|d|/j$, $X=K^{\varepsilon}$, and $Y=\sqrt{zuK|d|}/c\sqrt{j}$. It is straightforward to check that the conditions of the lemma are met under (9.3). We conclude that
@@
$\frac{-2\sqrt{2\pi}}{\sqrt{u}}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\chi_{d}(n)\sum_{(j,d)=1}\frac{1}{j}\textnormal{Im}\bigg(e^{-\pi i/2}\sum_{c=1}^{\infty}\frac{1}{[c,d]^{2}}\sum_{v\in\mathbb{Z}}\sum_{\eta\in\mathbb{Z}}\sum_{x\,([c,d])}\sum_{w\,([c,d])}\chi_{d}(x)\chi_{d}(w)$
(9.4) $\times S(xw,u;c)e\left(\frac{xv+w\eta}{[c,d]}\right)\int\limits_{\mathbb{R}}z^{-1}e\left(z\left(\frac{u[c,d]}{c^{2}v}-\frac{\eta}{[c,d]}\right)\right)W_{K}^{(2)}\left(\frac{zu[c,d]^{2}j}{v^{2}c^{2}|d|},\frac{zj}{|d|},\frac{K^{2}c^{2}v}{8\pi zu[c,d]}\right)\,\mathrm{d}z\bigg)$
$\qquad\qquad\qquad\qquad+O\left(X^{3/4}K^{1+\varepsilon}\right).$
@@
$v\ll_{\varepsilon}\frac{[c,d]|d|uK^{\varepsilon}}{c^{2}Kj}$
and the same holds for $\eta$. Note that as $|d|\ll X\ll\sqrt{K}$ and $u\leq K^{4\delta_{0}}\leq K^{\varepsilon}$ this shows that the integral is negligible unless $v,\eta\leq K^{2\varepsilon}/c$, which effectively truncates also the $c$-sum at $K^{2\varepsilon}$. Note also that we may truncate the $\eta$-sum at $X^{2}$ as otherwise the derivative of the phase function is $\gg[c,d]^{2}/c^{2}v$. Now using the first derivative test and estimating everything else trivially using the Weil bound for Kloosterman sums shows that (9) is $\ll X^{3/4}K^{1+\varepsilon}$.
@@
$-2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}c_{n}(u)\frac{\chi_{d}(nu)}{\sqrt{u}}$
@@
where the summand on the right-hand side is a multiplicative function of $c$. We also have by a simple calculation that, for $\textnormal{Re}(x+y)>1$,
$\sum_{c=1}^{\infty}\frac{\varphi(c)}{c^{1+x+y}}\cdot\frac{1}{d}\prod_{\begin{subarray}{c}p|d\\ p|c\end{subarray}}(-p)$ $=\frac{1}{d}\prod_{p}\left(1+\sum_{j=1}^{\infty}\frac{\varphi(p^{j})}{p^{j(1+x+y)}}\prod_{\begin{subarray}{c}q|d\\ q|p^{j}\end{subarray}}(-q)\right)$
@@
Making the change of variables $z\mapsto cz$, using the approximation (5.11) and Poisson summation we see that the integral in (9) equals
@@
$=\frac{1}{(2\pi i)^{2}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}\left(\frac{hcuj}{K|d|v^{2}}\right)^{-x}\left(\frac{hcj}{K|d|}\right)^{-y}\hbar_{x+y}\left(\frac{K^{2}|d|v}{8\pi hu}\right)\,\frac{\mathrm{d}x\,\mathrm{d}y}{xy}$
$=\frac{1}{(2\pi i)^{3}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(1-\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}\left(\frac{hcuj}{K|d|v^{2}}\right)^{-x}\left(\frac{hcj}{K|d|}\right)^{-y}\left(\frac{8\pi hu}{K^{2}|d|v}\right)^{z}\widetilde{\hbar}_{x+y}(z)\,\mathrm{d}z\,\frac{\mathrm{d}x\,\mathrm{d}y}{xy},$
@@
$2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)$
$\qquad\times\sum_{v|u}\frac{1}{(2\pi i)^{3}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(1-\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\widetilde{\hbar}_{x+y}(z)$
@@
$=2\sqrt{2\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}$
$\qquad\times\frac{1}{(2\pi i)^{3}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(1-\varepsilon)}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\widetilde{\hbar}_{x+y}(z)$
@@
$2K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}$
$\qquad\times\frac{1}{(2\pi i)^{3}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(1-\varepsilon)}\int\limits_{0}^{\infty}(2\pi)^{-x-y}e^{x^{2}+y^{2}}u^{z-x}(8\pi)^{z}K^{x+y-2z}d^{x+y-z}v^{2x-z}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\xi^{(x+y)/2-z}$
@@
$2K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{(2\pi i)^{2}}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{(1-\varepsilon)}\int\limits_{0}^{\infty}(2\pi K)^{-z}u^{z-x}(8\pi)^{z}e^{x^{2}+(z-x)^{2}}v^{2x-z}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\xi^{-z/2}$
@@
(9.7) $-2\zeta(0)K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)(-1)^{\#\{p|d\}}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}\int\limits_{0}^{\infty}e^{2x^{2}}u^{-x}v^{2x}\frac{h(\sqrt{\xi})}{\sqrt{\xi}}\prod_{p|d}\left(1-p^{-1}\right)\left(\frac{1-p}{p-1}\right)\,\mathrm{d}\xi\frac{\mathrm{d}x}{x^{2}}.$
Simplifying this by recalling that $\zeta(0)=-1/2$ and making a change of variables in the $\xi$-integral we see that (9) equals, up to a negligible error,
$2\widehat{h}(0)K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\frac{\varphi(d)}{d}\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\sum_{v|u}\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}e^{2x^{2}}u^{-x}v^{2x}\,\frac{\mathrm{d}x}{x^{2}}.$
@@
$M_{2}(X;d)$ $:=(\log K)\sum_{n\leq K^{4\delta_{0}}}\frac{h_{2}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u)$
@@
$=\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathbb{E}\left(L(X;d,k)M_{2}(X;d)\right)$
(9.8) $+2K(\log K)\widehat{h}(0)\sum_{n\leq K^{4\delta_{0}}}\frac{h_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\frac{\varphi(d)}{d}\phi\left(\frac{|d|}{X}\right)\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}\frac{e^{2x^{2}}}{x^{2}}\cdot u^{-x}\sum_{v|u}v^{2x}\,\mathrm{d}x$
$\qquad\qquad\qquad\qquad\qquad\qquad+O\left(X^{3/4}K^{1+\varepsilon}\right)$